please help: solve for v: -1/2=2/5v+2/3 :: (-1/2 the whole fraction is negative).
Answer by Alan3354(69443)

solve for v: -1/2=2/5v+2/3
I'll assume it's (2/5)v and not 2/(5v) or 2/(5v + 2/3)
Use parentheses, they're free.
solve for v: -1/2=(2/5)v + (2/3)
Multiply thru by 30
-15 = 12v + 20
12v = -35
v = -35/12